# Looking at Model Debiasing through the Lens of Anomaly Detection

### v2

@@ Abstract
It is widely recognized that deep neural networks are sensitive to bias in the data. This means that during training these models are likely to learn spurious correlations between data and labels, resulting in limited generalization abilities and low performance. In this context, model debiasing approaches can be devised aiming at reducing the model’s dependency on such unwanted correlations, either leveraging the knowledge of bias information or not. In this work, we focus on the latter and more realistic scenario, showing the importance of accurately predicting the bias-conflicting and bias-aligned samples to obtain compelling performance in bias mitigation. On this ground, we propose to conceive the problem of model bias from an out-of-distribution perspective, introducing a new bias identification method based on anomaly detection. We claim that when data is mostly biased, bias-conflicting samples can be regarded as outliers with respect to the bias-aligned distribution in the feature space of a biased model, thus allowing for precisely detecting them with an anomaly detection method. Coupling the proposed bias identification approach with bias-conflicting data upsampling and augmentation in a two-step strategy, we reach state-of-the-art performance on synthetic and real benchmark datasets. Ultimately, our proposed approach shows that the data bias issue does not necessarily require complex debiasing methods, given that an accurate bias identification procedure is defined.
@@ 1 Introduction
It is commonly acknowledged that deep neural networks can reach outstanding performance when used to tackle several computer vision tasks, including image classification, segmentation, and object detection, to name a few. Nonetheless, their performance is severely affected by the presence of bias in the data. In this case, deep neural networks often rely on spurious correlations between data and labels rather than capturing the semantic features of the objects of interest, leading to poor generalization performance. In this context, model debiasing methods have been adopted to reduce the model dependency on the spurious correlations in a biased dataset, enhancing the model capacity to capture semantic target features, while mitigating the bias effect. As an example, let us consider the Waterbirds dataset associated with the binary classification task of predicting images of water and land birds [1]. In this collection, most images of water birds show water in the background, which is generally absent in images of land birds. Consequently, a deep neural network trained on such data learns an undesired shortcut, likely focusing on the background to make predictions, rather than the actual appearance of the birds.
@@
In the last few years, several methods have been proposed to deal with biased datasets. The existing works for bias mitigation either exploit ground-truth bias annotations (regarded as supervised debiasing methods) or do not assume any prior knowledge of the bias (denoted as unsupervised debiasing methods). While supervised approaches operate a finer control over the learning process allowing to selectively mitigate the bias effects, and typically leading to higher performances, unsupervised methods operate blindly, both bias attributes and biased samples are unknown, which is typically the case in a realistic scenario. Among the unsupervised debiasing methods, common approaches leverage 2-stage procedures aiming at, first, detecting the samples affected by the bias (denoted as bias-aligned) and those that are not (bias-conflicting); and second, applying an actual debiasing process exploiting the estimated (and approximate) split of the data [2, 3, 4, 5, 6]. Empirical evidence shows that the more accurate the prediction of bias-conflicting samples, the higher the generalization performances of the debiased model. This is key to the entire process, yet quite challenging since prior (bias) information is missing, but pervades the large majority of the data, making extremely problematic the detection of the few bias-conflicting samples.
@@
On this ground, we propose that bias-conflicting samples can be considered as outliers in a biased distribution of the feature space inferred by a (biased) model trained on this data. To support our intuition, we train a ResNet-18 model for classification on the training set of Corrupted CIFAR-10, a synthetic biased dataset typically used as a benchmark for model debiasing. We then extract the embeddings of the training set, on which we apply Principal Component Analysis (PCA) for visualization. Then, we input the trained model with bias-conflicting and bias-aligned test samples to extract their embeddings and project them using the same Principal Component (PC) transformation applied to the training set. In Fig. 1, we show the first two PCs of the feature space obtained in this way (green dots for bias-aligned and red dots for bias-conflicting in the figure). For the sake of simplicity, we visualize the feature space for the Cat class only. From this figure, one can notice that bias-conflicting samples experience a shift in the feature space distribution of the ResNet-18 model, as compared to the distribution of the bias-aligned ones. This simple example evidences the existence of a distribution shift between bias-conflicting and bias-aligned samples and, given the strong imbalance of such subsets, suggests the usage of an anomaly detection method. Specifically, the shift is limited to the marginal distribution of the embeddings, it is present in each class (in different forms depending on the type of bias), and it can be attributed to a covariate shift in the data [7]. This can be cast as an outlier detection problem, for which anomaly detection methods can be successfully adopted to estimate test samples with covariate shifts [8]. Hence, we propose that an accurate bias-conflicting sample detection can be obtained regarding such samples as anomalies in the embedded feature space of an intentionally biased model. This is in fact a model specifically trained to learn bias-aligned samples, discouraged from learning bias-conflicting samples, and thus amplifying the shift in distribution from the bias-aligned ones.
@@
To sum up, the main contribution of this work is as follows. (i) We claim that bias-conflicting samples generate anomalous feature embeddings in the feature space of a biased model. Thus, we propose the usage of an anomaly detection method for identifying bias-conflicting (and aligned) samples, leveraging an intentionally biased model. To the best of our knowledge, this is the first work that proposes to look at bias-conflicting samples as anomalies, paving the way to the usage of anomaly detection methods for model debiasing tasks, and potentially opening new scenarios to bridge the gap between these two research areas. (ii) We validate our proposed bias-identification approach coupling it with a debiasing method based on predicted bias-conflicting upsampling and data augmentation to fine-tune a model pre-trained on bias data with vanilla ERM. We refer to the resulting two-step method as MoDAD, Model Debiasing with Anomaly Detection. Our results show that our approach based on anomaly detection bias identification reaches competitive performances on a synthetic dataset and state-of-the-art accuracy for three real-world benchmark datasets, obtaining an average improvement of $\sim 3\%$.
@@ 2 Related Works
The literature addressing data bias and model debiasing is rich, presenting diverse approaches and methodologies that aim at reducing the impact of bias on the generalization capabilities of neural networks. Among them, supervised debiasing methods utilize explicit bias annotations, while unsupervised approaches do not rely on any prior information about the bias. Existing supervised debiasing methods include distributionally robust optimization [1], and bias feature disentanglement [9]. In ReBias, Bahng et al[10] propose an approach relying on prior knowledge on the bias rather than bias labels. Wang et al[11] introduce HEX, a method based on a differentiable neural building block, to reduce dependence on image textural information.
@@
We focus on state-of-the-art analysis on unsupervised debiasing approaches, which share the same high-level structure of our proposed method. Nam et al[12] propose an architecture based on two models. The first model is intentionally biased using a generalized cross-entropy loss function, while a second model is jointly trained with the first one with a weighted cross-entropy loss function, under the assumption that bias-conflicting samples would contribute more to the loss than the bias-aligned ones. Lee et al[13] demonstrate the importance of training a model with diverse bias-conflicting samples, proving that “diversity” outperforms oversampling. They synthesize diverse bias-conflicting samples from the bias-aligned ones.
@@
Among the unsupervised methods, a common approach involves the usage of two-step algorithms, where the first step consists in estimating the correct split between bias-conflicting and bias-aligned samples, which is further exploited for debiasing the model, typically devoting more importance to bias-conflicting samples during training. In [2], a model with an early stopping criterion based on explicit bias information in a validation set is first trained. The model’s misclassified samples are then selected as bias-conflicting, and a second model is trained with a weighted cross-entropy loss function, whose weights are set proportionally to the number of identified bias-aligned and conflicting samples. Kim et al[4] provide a data augmentation framework to generate “bias-swapped” images, accounting for the bias attributes from the bias-conflicting images, and the bias-irrelevant attributes from the bias-aligned samples. Their proposed bias identification strategy consists in assigning a bias score to each sample based on both the classification correctness and model confidence, using a model trained with Generalized Cross-Entropy (GCE) loss. Seo et al[14] propose a method based on feature clustering and cluster re-weighting, starting from the observation that examples with the same bias attributes tend to have similar representations in the feature space. In [5], assuming that the training samples are partitioned into domains or environments, the authors aim to find environments that maximally violate the invariant learning principle. In these environments, the reference classifier associates the same feature vector with examples of different classes. Features that are domain-invariant predict the true class regardless of the domain. A similar bias identification approach is used in [6]. In [3], the hidden stratification phenomenon is tackled, which happens when classes comprise multiple finer-grained sub-classes. The identification of the sub-classes is achieved through feature clustering applied on a model trained on the super-class classification task.
@@ 3.2.1 OCSVM customization
Such decision function is outlined in Eq. (1), where $\alpha_{i}$ and $\lambda$ are the learned Lagrange multipliers and hyperplane’s offset respectively, while $K$ is the Gram matrix of distances between sample points in kernel space. The higher the argument of the sign, the higher the (signed) distance of a sample from the learned margin, with a positive distance indicating that a sample should be regarded as in-class (positive sign), or out-of-class (negative sign). As such, the argument of the sign, can be regarded as an anomaly score.
In our bias identification approach, we modify Eq. (1) to shift the argument of the sign by a threshold $\tau$ (which we automatically compute, see Algorithm 1), so that when the anomaly score is higher than $\tau$ the corresponding sample is classified as in-class, and anomalous otherwise. Such a decision threshold can be used to restrict anomaly predictions with respect to the original formulation. In our approach, we train one OCSVM per target class, allowing us to tailor bias detection specifically to individual categories.
@@ 3.2.2 Amplifying the distribution shift with GCE
5:  Extract training samples embeddings $\mathcal{Z}=\{\left(\Phi(x),y\right),\;\forall\>(x,y)\in\mathcal{D}\}$.
@@
8:  Compute a decision threshold $\tau_{y}$ as the anomaly score corresponding to $\textnormal{Percentile}(\mathcal{A}_{y},\alpha_{y})$, $\alpha_{y}=100\cdot\frac{1}{2}\left(\frac{\psi_{y}-|\mathcal{C}_{y}|}{\psi_{y}}\right)$
@@ 3.3 Bias Mitigation: Fine-Tuning the Biased Model
In this work, we design a debiasing scheme based on bias-conflicting upsampling and augmentation (see Fig. 2). Starting from a model trained using the biased dataset with vanilla ERM and hence likely biased, we employ a weighted random sampler with replacement, whose weights are set depending on the prediction $\hat{\mathcal{B}}$ of the bias identification step, as the inverse of the two estimated populations.
With this sampler, during the debiasing training process, we build each mini-batch such that the ratio of bias-aligned and bias-conflicting samples is roughly balanced. We further perform upsampling adding three augmented images for each bias-conflicting sample in each mini-batch, so that the number of bias-conflicting samples in the batch results 4 times higher than the number of bias-aligned ones. Exploiting this batch construction scheme, a classic CE loss function is then used for debiasing a model. Differently from other two-step methods (e.g., JTT [2]), MoDAD performs the debiasing process on an already biased model, facing a realistic scenario when bias is unknown.
@@ 4.1 Datasets
Figure 3: Example of bias-aligned (left) and bias-conflicting samples (right) from Corrupted CIFAR-10 Dataset.
@@
Figure 5: Example of bias-aligned (left) and bias-conflicting samples (right) from BFFHQ dataset.
Waterbirds is an image dataset introduced in [1] that combines bird subjects from the Caltech-UCSD Birds-200-2011 (CUB) dataset [24] and image backgrounds from the Places dataset [25]. Images are divided according to a target attribute regarding birds being Waterbirds or Landbirds, while the background being land or water represents the bias attribute. In this setting, images depicting Landbirds on land and Waterbirds on water are bias-aligned samples, whereas pairing a Landbird with a water background constitutes a bias-conflicting sample (and vice-versa for Waterbirds on land background).
Figure 6: Example of bias-aligned (left) and bias-conflicting samples (right) from Waterbirds dataset.
@@ 4.3 Results
Table 1: MoDAD’s Average Accuracy on Corrupted CIFAR-10 and related comparisons. BS refers to Bias Supervision. Best results in bold.
Table 1 provides MoDAD’s results on three different $\rho$ values of Corrupted CIFAR-10, benchmarking with relevant existing approaches. Despite the basic debiasing step, our method is competitive with respect to state-of-the-art approaches for the three different bias ratios, reaching similar performances of DFA [13] and LfF [12], while being slightly lower for $\rho=0.995$. Please, note that our performance result is more stable considering the standard deviation, which is an order of magnitude lower than that of the other methods. Moreover, we significantly outperform JTT [2], whose debiasing approach is similar to the one employed in MoDAD, for all the bias ratios, further suggesting the effectiveness of our bias identification mechanism. Regarding the accuracy of our anomaly detection algorithm in bias identification, the average and standard deviation for the F1-score over the ten corrupted CIFAR-10 classes, correspond to $0.81\pm 0.05$, $0.80\pm 0.07$ and $0.68\pm 0.11$, for $\rho=\{0.950,0.980,0.995\}$ respectively (see Table 2). Finally, in the last entry of Table 1 we report the oracle represented by using the ground truth bias labels and our proposed debiasing method. The reported oracle accuracy provides the best performances among the referenced methods, confirming our intuition regarding the importance of precise bias identification for model debiasing.
Dataset $\bf\rho$ F1-Score
Corrupted CIFAR-10 $0.950$ $0.81\pm 0.05$
Corrupted CIFAR-10 $0.980$ $0.80\pm 0.07$
Corrupted CIFAR-10 $0.995$ $0.68\pm 0.11$
Waterbirds $0.950$ $0.70\pm 0.07$
BFFHQ $0.995$ $0.75\pm 0.05$
Table 2: F1-Score of our per-class Bias Identification method (standard deviation refers to the target classes)
@@
Average Conflicting Average Conflicting Average Conflicting

### v3

@@ Abstract
Deep neural networks are likely to learn unintended spurious correlations between training data and labels when dealing with biased data, potentially limiting the generalization to unseen samples not presenting the same bias. In this context, model debiasing approaches can be devised aiming at reducing the model’s dependency on such unwanted correlations, either leveraging the knowledge of bias information or not. In this work, we focus on the latter and more realistic scenario, showing the importance of accurately predicting the bias-conflicting and bias-aligned samples to obtain compelling performance in bias mitigation. On this ground, we propose to conceive the problem of model bias from an out-of-distribution perspective, introducing a new bias identification method based on anomaly detection. We claim that when data is mostly biased, bias-conflicting samples can be regarded as outliers with respect to the bias-aligned distribution in the feature space of a biased model, thus allowing for precisely detecting them with an anomaly detection method. Coupling the proposed bias identification approach with bias-conflicting data upsampling and augmentation in a two-step strategy, we reach state-of-the-art performance on synthetic and real benchmark datasets. Ultimately, our proposed approach shows that the data bias issue does not necessarily require complex debiasing methods, given that an accurate bias identification procedure is defined. Source code is available at https://github.com/Malga-Vision/MoDAD
footnotetext: Correspondence: vito.paolo.pastore@unige.itfootnotetext: Work accepted at IEEE/CVF Winter Conference on Applications of Computer Vision (WACV) 2025.
@@ 1 Introduction
It is commonly acknowledged that deep neural networks can reach outstanding performance when used to tackle several computer vision tasks, including image classification, segmentation, and object detection, to name a few. Nonetheless, in case of biased data, deep neural networks often rely on spurious correlations between data and labels rather than capturing the semantic features of the objects of interest, leading to poor generalization performance. In this context, model debiasing methods have been adopted to reduce the model dependency on the spurious correlations in a biased dataset, enhancing the model capacity to capture semantic target features, while mitigating the bias effect. As an example, let us consider the Waterbirds dataset associated with the binary classification task of predicting images of water and land birds [30]. In this collection, most images of water birds show water in the background, which is generally absent in images of land birds. Consequently, a deep neural network trained on such data learns an undesired shortcut, likely focusing on the background to make predictions, rather than the actual appearance of the birds.
@@
In the last few years, several methods have been proposed to deal with biased datasets. The existing works for bias mitigation either exploit ground-truth bias annotations (regarded as supervised debiasing methods) or do not assume any prior knowledge of the bias (denoted as unsupervised debiasing methods). While supervised approaches operate a finer control over the learning process allowing to selectively mitigate the bias effects, and typically leading to higher performances, unsupervised methods operate blindly, both bias attributes and biased samples are unknown, which is typically the case in a realistic scenario.
Among the unsupervised debiasing methods, common approaches leverage 2-stage procedures aiming at, first, detecting the samples affected by the bias (denoted as bias-aligned) and those that are not (bias-conflicting); and second, applying an actual debiasing process exploiting the estimated (and approximate) split of the data [19, 33, 14, 5, 1]. Among existing debiasing works, the general assumption is that bias-aligned represents the majority of training samples (from $95\%$ to $99.5\%$ in common benchmark datasets), and we share this setting in our work. Empirical evidence shows that the more accurate the prediction of bias-conflicting samples, the higher the generalization performances of the debiased model. This is key to the entire process, yet quite challenging since prior (bias) information is missing, but pervades the large majority of the data, making extremely problematic the detection of the few bias-conflicting samples.
@@
On this ground, we propose that bias-conflicting samples can be considered as outliers in a biased distribution of the feature space inferred by a (biased) model trained on this data. To support our intuition, we provide an example in Fig. 1. Specifically, we train a ResNet-18 model for classification on the training set of Corrupted CIFAR-10, a synthetic biased dataset typically used as a benchmark for model debiasing. We then extract the embeddings of the training set, on which we apply Principal Component Analysis (PCA) for visualization. Then, we input the trained model with bias-conflicting and bias-aligned test samples to extract their embeddings and project them using the same Principal Component (PC) transformation applied to the training set. In Fig. 1, we show the first two PCs of the feature space obtained in this way (green dots for bias-aligned and red dots for bias-conflicting in the figure). For the sake of simplicity, we visualize the feature space for the Cat class only. From this figure, one can notice that bias-conflicting samples experience a shift in the feature space distribution of the ResNet-18 model, as compared to the distribution of the bias-aligned ones. This simple example evidences the existence of a distribution shift between bias-conflicting and bias-aligned samples and, given the strong imbalance of such subsets, suggests the usage of an anomaly detection method. Specifically, the shift is limited to the marginal distribution of the embeddings, it is present in each class (in different forms depending on the type of bias), and it can be attributed to a covariate shift in the data [24]. This can be cast as an outlier detection problem, for which anomaly detection methods can be successfully adopted to estimate test samples with covariate shifts [42].
Hence, we propose that an accurate bias-conflicting sample detection can be obtained regarding such samples as anomalies in the embedded feature space of an intentionally biased model. This is in fact a model specifically trained to learn bias-aligned samples, discouraged from learning bias-conflicting samples, and thus amplifying the shift in distribution from the bias-aligned ones. Furthermore, we cast the bias identification as a per-class problem and exploit one anomaly detector per target class. This is a fundamental aspect of our method, as the distribution shift is amplified in the feature space of single classes, thus allowing an accurate bias-conflicting identification.
@@
To sum up, the main contribution of this work is as follows. (i) We claim that bias-conflicting samples generate anomalous feature embeddings in the feature space of a biased model. Thus, we propose the usage of an anomaly detection method for identifying bias-conflicting (and aligned) samples per target class, leveraging an intentionally biased model. To the best of our knowledge, this is the first work that proposes to look at bias-conflicting samples as anomalies, paving the way to the usage of anomaly detection methods for model debiasing tasks, and potentially opening new scenarios to bridge the gap between these two research areas. (ii) We validate our proposed bias-identification approach coupling it with a debiasing method based on predicted bias-conflicting upsampling and data augmentation to fine-tune a model pre-trained on bias data with vanilla ERM. We refer to the resulting two-step method as MoDAD, Model Debiasing with Anomaly Detection. Our results show that our approach based on anomaly detection bias identification reaches competitive performances on a synthetic dataset and state-of-the-art accuracy for three real-world benchmark datasets, obtaining an average improvement of $\sim 3\%$.
@@ 2 Related Works
The literature addressing data bias and model debiasing is rich, presenting diverse approaches and methodologies that aim at reducing the impact of bias on the generalization capabilities of neural networks. Among them, supervised debiasing methods utilize explicit bias annotations, while unsupervised approaches do not rely on any prior information about the bias. Existing supervised debiasing methods include distributionally robust optimization [30], and bias feature disentanglement [35]. In ReBias, Bahng et al. [3] propose an approach relying on prior knowledge on the bias rather than bias labels. Wang et al. [41] introduce HEX, a method based on a differentiable neural building block, to reduce dependence on image textural information.
@@
We focus on state-of-the-art analysis on unsupervised debiasing approaches, which share the same high-level structure of our proposed method. Nam et al. [25] propose an architecture based on two models. The first model is intentionally biased using a generalized cross-entropy loss function, while a second model is jointly trained with the first one with a weighted cross-entropy loss function, under the assumption that bias-conflicting samples would contribute more to the loss than the bias-aligned ones. Lee et al. [17] demonstrate the importance of training a model with diverse bias-conflicting samples, proving that “diversity” outperforms oversampling. They synthesize diverse bias-conflicting samples from the bias-aligned ones.
@@
Among the unsupervised methods, a common approach involves the usage of two-step algorithms, where the first step consists in estimating the correct split between bias-conflicting and bias-aligned samples, which is further exploited for debiasing the model, typically devoting more importance to bias-conflicting samples during training. In [19], a model with an early stopping criterion based on explicit bias information in a validation set is first trained. The model’s misclassified samples are then selected as bias-conflicting, and a second model is trained with a weighted cross-entropy loss function, whose weights are set proportionally to the number of identified bias-aligned and conflicting samples. Kim et al. [14] provide a data augmentation framework to generate “bias-swapped” images, accounting for the bias attributes from the bias-conflicting images, and the bias-irrelevant attributes from the bias-aligned samples. Their proposed bias identification strategy consists in assigning a bias score to each sample based on both the classification correctness and model confidence, using a model trained with Generalized Cross-Entropy (GCE) loss. Seo et al. [32] propose a method based on feature clustering and cluster re-weighting, starting from the observation that examples with the same bias attributes tend to have similar representations in the feature space. In [5], assuming that the training samples are partitioned into domains or environments, the authors aim to find environments that maximally violate the invariant learning principle. In these environments, the reference classifier associates the same feature vector with examples of different classes. Features that are domain-invariant predict the true class regardless of the domain. A similar bias identification approach is used in [1]. In [33], the hidden stratification phenomenon is tackled, which happens when classes comprise multiple finer-grained sub-classes. The identification of the sub-classes is achieved through feature clustering applied on a model trained on the super-class classification task.
@@ 3.2.1 OCSVM customization
Such decision function is outlined in Eq. (1), where $\alpha_{i}$ and $\lambda$ are the learned Lagrange multipliers and hyperplane’s offset respectively, while $K$ is the Gram matrix of distances between sample points in kernel space. The higher the argument of the sign, the higher the (signed) distance of a sample from the learned margin, with a positive distance indicating that a sample should be regarded as in-class (positive sign), or out-of-class (negative sign). As such, the argument of the sign, can be regarded as an anomaly score. In our bias identification approach, we modify Eq. (1) to shift the argument of the sign by a threshold $\tau$ (which we automatically compute, see Algorithm 1), so that when the anomaly score is higher than $\tau$ the corresponding sample is classified as in-class, and anomalous otherwise. Such a decision threshold can be used to restrict anomaly predictions with respect to the original formulation.
@@ 3.2.2 Amplifying the distribution shift with GCE
5:  Extract training samples embeddings and respective labels in a set $\mathcal{Z}=\{\left(\Phi(x),y\right),\;\forall\>(x,y)\in\mathcal{D}\}$.
@@
8:  Compute a decision threshold $\tau_{y}$ as the anomaly score corresponding to $\textnormal{Percentile}(\mathcal{A}_{y},\pi_{y})$, $\pi_{y}=100\cdot\frac{1}{2}\left(\frac{\psi_{y}-|\mathcal{C}_{y}|}{\psi_{y}}\right)$
@@
3.2.3 Identifying conflicting samples at a class level
In our approach, we train an OCSVM per target class, allowing us to tailor bias detection specifically to individual categories. This is a core design choice of our approach, as we assume that the most dramatic shift happens inside the feature space of an individual class. As we employ GCE loss to train the biased model of our first step, we can expect that it will produce compact features for bias-aligned samples [25, 14, 17], while spreading around bias-conflicting features and consistently misclassifying a great portion of them. In addition to mitigating the memorization effect, this also results in an amplified separation between bias-conflicting and bias-aligned features of a specific class: conflicting features will mainly overlap with features from other classes (as they are misclassified) and thus greatly differ from the normality of its bias-aligned samples. This enables our customized class-specific anomaly detectors to better learn the distinction between samples with the same semantics but highly differing representations.
@@ 3.3 Bias mitigation: fine-tuning the biased model
In this work, we design a debiasing scheme based on bias-conflicting upsampling and augmentation (see Fig. 2). Starting from a model trained using the biased dataset with vanilla ERM and hence likely biased, we employ a weighted random sampler with replacement, whose weights are set depending on the prediction $\hat{\mathcal{B}}$ of the bias identification step, as the inverse of the two estimated populations. With this sampler, during the debiasing training process, we build each mini-batch such that the ratio of bias-aligned and bias-conflicting samples is roughly balanced. We further perform upsampling adding three augmented images for each bias-conflicting sample in each mini-batch, so that the number of bias-conflicting samples in the batch results 4 times higher than the number of bias-aligned ones. Exploiting this batch construction scheme, a classic CE loss function is then used for debiasing a model. Differently from other two-step methods (e.g., JTT [19]), MoDAD performs the debiasing process on an already biased model, facing a realistic scenario when bias is unknown.
@@ 4.1 Datasets
Figure 3: Example of bias-aligned (top) and bias-conflicting samples (bottom) from Corrupted CIFAR-10 Dataset.
@@
Figure 5: Example of bias-aligned (top) and bias-conflicting samples (bottom) from BFFHQ dataset.
@@
Waterbirds is an image dataset introduced in [30] that combines bird subjects from the Caltech-UCSD Birds-200-2011 (CUB) dataset [40] and image backgrounds from the Places dataset [44]. Images are divided according to a target attribute regarding birds being Waterbirds or Landbirds, while the background being land or water represents the bias attribute. In this setting, images depicting Landbirds on land and Waterbirds on water are bias-aligned samples, whereas pairing a Landbird with a water background constitutes a bias-conflicting sample (and vice-versa for Waterbirds on land background). Samples of this dataset are shown in Figure 6.
Figure 6: Example of bias-aligned (top) and bias-conflicting samples (bottom) from Waterbirds dataset.
@@ 4.3 Results
Table 2 provides MoDAD’s results on three different $\rho$ values of Corrupted CIFAR-10, benchmarking with relevant existing approaches. Despite the basic debiasing step, our method is competitive with respect to state-of-the-art approaches for the three different bias ratios, reaching similar performances of DFA [17] and LfF [25] while being slightly lower for $\rho=0.995$. Additionally, we significantly outperform JTT [19], whose debiasing approach is similar to the one employed in MoDAD, for all the bias ratios, further suggesting the effectiveness of our bias identification mechanism. Regarding the accuracy of our anomaly detection algorithm in bias identification, the average and standard deviation for the F1-score over the ten corrupted CIFAR-10 classes correspond to $0.81\pm 0.05$, $0.80\pm 0.07$ and $0.68\pm 0.11$, for $\rho=\{0.950,0.980,0.995\}$ respectively (see Table S.2). Finally, in the last entry of Table 2 we report the oracle represented by using the ground truth bias labels and our proposed debiasing method. The reported oracle accuracy provides the best performances among the referenced methods, confirming our intuition regarding the importance of precise bias identification for model debiasing.
Table 1: F1-Score of our per-class Bias Identification method (standard deviation refers to the target classes). We provide a comparison with other works reporting this metric (or precision and recall) on the explored datasets equipped with bias annotations.
Dataset $\bf\rho$ Ours BiaSwap GEORGE JTT
Corrupted CIFAR-10 $0.950$ $\bf 0.81\pm 0.05$ $0.72$
Corrupted CIFAR-10 $0.980$ $\bf 0.80\pm 0.07$ $0.69$
Corrupted CIFAR-10 $0.995$ $\bf 0.68\pm 0.11$ $0.63$
Waterbirds $0.950$ $\bf 0.70\pm 0.07$ $0.36$ $0.42$
BFFHQ $0.995$ $\bf 0.75\pm 0.05$ $0.68$
Table 2: MoDAD’s Average Accuracy on Corrupted CIFAR-10 and related comparisons. BS refers to Bias Supervision. Best results in bold, second best underlined.
@@
Average Average Conflicting Average Conflicting